# Five-loop Anomalous Dimensions of Cubic Scalar Theory from Operator Product Expansion

### v2

@@ 3.3 Anomalous dimensions of the $\varphi^{2}$ and $\varphi^{3}$ operators
The $\widetilde{Z}_{\varphi^{Q}}$-factors for the $\varphi^{Q}$ operator are determined by the UV finiteness conditions (29), and the corresponding two-point integrals are evaluated by (35). When $Q=2$, all contributing topologies are irreducible topologies. The tree-level, 1-loop and 2-loop topologies of $\varphi^{2}$-$2\varphi$ correlation function have been sketched in Fig.(4), together with their corresponding cut graphs.
@@
Note that due to the symmetry of graphs, a topology could generate more than one non-isomorphic cut graphs, similar as we have mentioned in the three-point function computation. For the cut graphs related to the same irreducible topology, the symmetry factors as well as the coupling factors are the same as those of the topology. Now Let us compute the $Z$-factor of $\varphi^{2}$, from eqn.(35) we get the explicit expression for $\varphi^{2}$ as
@@
There is only one 1-loop irreducible topology $\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi{\varphi}}$, which generates one cut graph $\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}}$. With results,
$\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=1~~~,~~~\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=g^{2}_{0}~~~,~~~\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{\epsilon}+2+\epsilon\left(\frac{5}{2}-\frac{\pi^{2}}{24}\right)+\mathcal{O}(\epsilon^{2})~,$ (65)
we get the required UV contribution of all 1-loop irreducible topologies for $\varphi^{2}\to 2\varphi$ as
@@
Note that in (65), an overall factor $(1/p^{2})^{\frac{\epsilon}{2}L}$ with $L=1$ is not explicitly shown in the result of UV divergence, while there will also be a factor $\mu^{\epsilon L}$ from the couplings $\mathbb{C}_{\footnotesize\mbox{f}}$ at $L$-loop. Hence for (66), if recovering the $(1/p^{2})^{\frac{\epsilon}{2}}$ factor and turning $g_{0},h_{0}\to g,h$, there will be an overall $\left(\frac{\mu^{2}}{p^{2}}\right)^{\frac{\epsilon}{2}}$ factor, which contributes to the $\ln\frac{\mu^{2}}{p^{2}}$ expansion. Generally, for $G_{\scriptsize\mbox{irre}}^{j{\tiny\mbox{-loop}}}$ there will be an overall $\left(\frac{\mu^{2}}{p^{2}}\right)^{j\frac{\epsilon}{2}}$ factor. For presentation purpose, the factor $(1/p^{2})^{\frac{\epsilon}{2}L}$ will always be treated as implicit in the expressions of UV divergences.
The 1-loop $\widetilde{Z}_{\varphi^{2}}$ can be determined from the UV finiteness of ${\widetilde{Z}_{\varphi^{2}}}G_{Q=2}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{Q})$, where we have used the fact that the 1PI contribution of $\widetilde{Z}_{\varphi^{1}}=1$. After rewriting all bare couplings $g_{0},h_{0}$ to $g,h$ by relations eqn.(7), then expanding the full expression at the limits $g,h\to 0$ followed by the limit $\epsilon\to 0$, we get the condition for the vanishing of $1/\epsilon$ pole in the leading Log term as,
$0=z_{\varphi^{2}}^{11}+g^{2}~~~\to~~~z_{\varphi^{2}}^{11}=-g^{2}~,$ (67)
and the sub-Log term provides no more conditions. Similarly, in order to determine the 2-loop $\widetilde{Z}_{\varphi^{2}}$, the contribution of $G_{\scriptsize\mbox{irre}}^{{\tiny\mbox{2-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{2})$ should be included. There are 5 independent irreducible topologies, which generates 7 cut graphs. The symmetry factors and coupling factors can be simply computed as,
@@
The UV divergences are again computed by TwoPoint command in HyperlogProcedures, which are
@@
Then the required UV contribution of all 2-loop irreducible topologies for $\varphi^{2}\to 2\varphi$ is
@@
After the series expansion of ${\widetilde{Z}_{\varphi^{2}}}G_{Q=2}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{Q})$, we get the UV finiteness conditions for the vanishing of $1/\epsilon$ and $1/\epsilon^{2}$ poles in the leading Log term as,
@@
The $z_{\varphi^{2}}^{11}$ has already been determined in the 1-loop computation, thus by solving above two equations we get,
@@
Note that the vanishing of $1/\epsilon$ pole in the sub-Log term provides an additional equation $0=g^{2}z_{\varphi^{2}}^{11}+g^{4}$. Hence even if we did not compute $z_{\varphi^{2}}^{11}$ at 1-loop, equations of leading Log and sub-Log terms are sufficient to solve $z_{\varphi^{2}}^{11},z_{\varphi^{2}}^{21},z_{\varphi^{2}}^{22}$, and completely determine the 2-loop $\widetilde{Z}_{\varphi^{2}}$-factor.
The determination of $\widetilde{Z}_{\varphi^{3}}$-factor follows exactly the same procedure, where now we should use the UV finiteness conditions of $\frac{\widetilde{Z}_{\varphi^{3}}}{\widetilde{Z}_{\varphi^{2}}}G_{Q=3}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},\mbox{\textst{$x$}}_{2},x_{Q})$. The topologies of $\varphi^{3}$-$3\varphi$ correlation function are sketched in Fig.(5).
@@
Notice that the loop topologies in the first row are equivalent to the $\varphi^{2}\to 2\varphi$ irreducible topologies we have already computed. There are only 2 new irreducible topologies $\Gamma^{\tiny\mbox{L2-XI}}_{\otimes\varphi{\varphi}\varphi}$, $\Gamma^{\tiny\mbox{L2-XII}}_{\otimes\varphi{\varphi}\varphi}$ at 2-loops for $\varphi^{3}\to 3\varphi$. Again from eqn.(35) we get the explicit expression for $\varphi^{3}$ as
@@
Since $G_{\scriptsize\mbox{irre}}^{{\tiny\mbox{1-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{2})$ and $G_{\scriptsize\mbox{irre}}^{{\tiny\mbox{2-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{2})$ have already been computed in $\varphi^{2}$ case as (66) and (77), in order to compute 2-loop $\widetilde{Z}_{\varphi^{3}}$-factor, we only need to compute the additional contributions from these 2 new irreducible topologies of $\varphi^{3}\to 3\varphi$. They generate in total 4 cut graphs as shown in Fig.(5). Their symmetry factors and coupling factors are given by
@@
Then the required UV contribution of all 2-loop irreducible topologies for $\varphi^{3}\to 3\varphi$ is
@@
With the previously computed $\widetilde{Z}_{\varphi^{2}}$-factor, we can take the series expansion for $\frac{\widetilde{Z}_{\varphi^{3}}}{\widetilde{Z}_{\varphi^{2}}}G_{Q=3}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},\mbox{\textst{$x$}}_{2},x_{Q})$, and get the following equations from the vanishing of $1/\epsilon$, $1/\epsilon^{2}$ poles in the leading Log term and $1/\epsilon$ pole in the sub-Log term,
@@
By solving these 3 equations we completely determine the 2-loop $\widetilde{Z}_{\varphi^{3}}$-factor as
@@
With results (67), (80) and (90), the 2-loop $\widetilde{Z}_{\varphi^{2}}$, $\widetilde{Z}_{\varphi^{3}}$-factors are finally determined as,
@@
The gamma functions of $\varphi^{Q}$ can be computed by $\gamma_{\varphi^{Q}}=-\frac{\partial\ln\widetilde{Z}_{\varphi^{Q}}}{\partial\ln\mu}$, hence we get
$\gamma_{\varphi^{2}}=-\frac{\partial\ln\widetilde{Z}_{\varphi^{2}}}{\partial g}\beta(g)-\frac{\partial\ln\widetilde{Z}_{\varphi^{2}}}{\partial h}\beta(h)=-g^{2}+\left(\frac{7g^{4}N}{72}-\frac{49g^{4}}{36}-\frac{g^{3}h}{2}+\frac{7g^{2}h^{2}}{72}\right)~,$ (92)
$\gamma_{\varphi^{3}}=-\frac{\partial\ln\widetilde{Z}_{\varphi^{3}}}{\partial g}\beta(g)-\frac{\partial\ln\widetilde{Z}_{\varphi^{3}}}{\partial h}\beta(h)=-3g^{2}+\left(\frac{7g^{4}N}{24}-\frac{67g^{4}}{12}-\frac{5g^{3}h}{2}+\frac{7g^{2}h^{2}}{24}\right)~,$ (93)
where $\beta(g),\beta(h)$ have been provided in (61), (62). By the ansatz that the 2-loop $\gamma_{\varphi^{Q}}$ takes the form,
$\gamma_{\varphi^{Q}}=c_{1}Q(Q-1)(Q-2)+c_{2}Q(Q-1)~,$ (94)
and from results $\gamma_{\varphi^{2}},\gamma_{\varphi^{3}}$ we determine
@@
Then the scaling dimension of the $\varphi^{Q}$ operator can be calculated by its definition as,
$\Delta_{Q}=\left(\frac{d}{2}-1\right)Q+\gamma_{\varphi^{Q}}+Q\gamma_{\varphi}~,$ (96)
where the anomalous dimension of the field $\varphi$ has been computed in (63).
@@ 4 Five-loop anomalous dimensions
The procedures described in previous section for 2-loop computations can be straightforwardly generalized to higher-loop calculations following exactly the same manner, with the primary difference being the need to consider a larger number of graphs. The $Z_{\varphi},Z_{\sigma}$-factors are determined from their corresponding two-point topologies in the manner analogous to (41). Similarly, the $Z_{h},Z_{g}$-factors are determined from the two-point integrals of cut graphs, generated by removing one external leg from the respective three-point topologies, following the approach outlined in (55). The beta functions are then computed from $\widetilde{Z}_{g},\widetilde{Z}_{h}$ using the definition (9). The field anomalous dimensions $\gamma_{\varphi}$ and $\gamma_{\sigma}$ can be computed from the $Z_{\varphi},Z_{\sigma}$-factors combined with the beta functions, following the definition (10). The $\widetilde{Z}_{\varphi^{Q}}$ factor is determined from the two-point integrals of cut graphs generated by removing (Q1)(Q-1) external legs from irreducible 1PI $\varphi^{Q}\to Q\varphi$ topologies, in a manner similar to that described in eqn.(35). The anomalous dimension of the $\varphi^{Q}$ operator is then given by $\gamma_{\varphi^{Q}}+Q\gamma_{\varphi}$, where $\gamma_{\varphi^{Q}}$ is computed from $\widetilde{Z}_{\varphi^{Q}}$ and the beta functions by its definition. Determining the 5-loop $\gamma_{\varphi^{Q}}$ for general $Q$ requires the results of $L$-loop $\gamma_{\varphi^{Q_{L}}}$ with $Q_{L}=2,\ldots,L+1$ up to 5-loops. Since UV divergences are computed from scalar integrals, we do not need to consider the specific field configurations of Feynman diagrams, but only their topologies.
All topologies for two-point, three-point and the $\varphi^{Q}$-$Q\varphi$ correlation functions, as well as their cut graphs are generated by private Mathematica codes. The number of 1PI topologies for two-point and three-point functions, and irreducible correlation functions is listed in Table.(2).
@@
The number grows rapidly as the number of loops increases. Cut graphs are generated through non-isomorphic computations of all possible external leg removals, and each of these cut graphs should be computed to determine their UV divergences. The number of cut graphs is listed in Table.(2). It can be seen that the number of cut graphs to be computed is about 3 to 4 times larger than the number of the original topologies.
The UV divergences are computed via TwoPoint command in HyperlogProcedures Maple package. The computational capability and time consumption of these evaluations are sensitive to both the loop order and the order of the dimensional regulator $\epsilon$. To apply the UV finiteness conditions for $L_{\tiny\mbox{max}}$-loop computation, an $L$-loop divergence in dimensional regularization should be computed up to the $\epsilon^{L_{\tiny\mbox{max}}-1-L}$ order. Increasing $L_{\tiny\mbox{max}}$ by one order requires UV divergence computations with an additional order in $\epsilon$ for all loop integrals, which significantly increases time consumption. For 5-loop computations, an $L$-loop divergence should be computed to the $\epsilon^{4-L}$ order, and the current version of HyperlogProcedures is capable of evaluating all required two-point integrals. Time consumption varies depending on the computer’s CPU and memory usage. On a normal desktop with a CPU base frequency 3.20GHz and 64GB memory, the approximate time required to evaluate all loop integrals is listed in Table.(3).
@@
Time consumption also increases rapidly as the loop order rises. However, for 5-loop computations, it remains within a reasonably acceptable range. In contrast, the symmetry factors and coupling factors for these topologies are generated almost instantaneously compared to the time required for evaluating UV divergences.
We have computed $Z_{\varphi}$, $Z_{\sigma}$, $Z_{h}$, $Z_{g}$ and all $\widetilde{Z}_{\varphi^{Q_{L}}},Q_{L}=2,\ldots,L+1$ up to 5-loops. From these results, we have derived the beta functions $\beta(g),\beta(h)$, the field anomalous dimensions $\gamma_{\varphi}$, $\gamma_{\sigma}$, and the scaling dimensions of the $\varphi^{Q}$} operator up to 5-loops. Results of lower than 5-loops have been compared with existing literature with agreements Fei:2014xta ; Gracey:2015tta ; Kompaniets:2021hwg .
The 5-loop correction to the anomalous dimension $\gamma_{\varphi^{Q}}+Q\gamma_{\varphi}$ is presented in Appendix §A. The scaling dimension of the $\varphi^{Q}$ operator at the Wilson-Fisher fixed point can be computed as
$\Delta_{Q\tiny\mbox{FP}}=\left(2-\frac{\epsilon}{2}\right)Q+\left(\gamma_{\varphi^{Q}}+Q\gamma_{\varphi}\right)\Big{|}_{g=g^{\ast},h=h^{\ast}}~,$ (97)
where $g^{\ast},h^{\ast}$ are the fixed-point solutions of the beta functions. Our 5-loop fixed-point solutions, provided in Appendix §B, are in perfect agreement with the known results in the literature Kompaniets:2021hwg . By substituting these fixed-point solutions into the anomalous dimensions, we obtain the $\Delta_{Q}$ in the large $N$ limit. The leading, sub-leading and sub-sub-leading order contributions of the large $N$ expansion are as follows,

### v4

@@ 3.3 Anomalous dimensions of the $\varphi^{2}$ and $\varphi^{3}$ operators
The $\widetilde{Z}_{\varphi^{Q}}$-factors for the $\varphi^{Q}$ operator are non-trivially related to two-point integrals via UV finiteness conditions in eqn.(29), where the two-point integrals are generated by removing (Q1)(Q-1) external legs from $\varphi^{Q}\to Q\varphi$ correlation function graphs. These integrals are evaluated using eqn.(35). For $Q=2$, all contributing topologies are irreducible. The tree-level, one-loop and two-loop topologies of the $\varphi^{2}$-$2\varphi$ correlation function, along with their corresponding cut graphs, are illustrated in Fig.(4).
@@
Due to graph symmetry, a single topology may generate multiple non-isomorphic cut graphs (as observed in three-point function computations), and cut graphs derived from the same irreducible topology share identical symmetry and coupling factors. For $Q=2$, eqn.(35) yields the explicit expression for $\varphi^{2}$,
@@
There is only one 1-loop irreducible topology $\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi{\varphi}}$, which generates the cut graph $\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}}$ with the following inputs,
$\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=1~~~,~~~\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=g^{2}_{0}~~~,~~~\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{\epsilon}+2+\epsilon\left(\frac{5}{2}-\frac{\pi^{2}}{24}\right)+\mathcal{O}(\epsilon^{2})~.$ (65)
The one-loop irreducible UV contribution for $\varphi^{2}\to 2\varphi$ is
@@
Note that the $L=1$ loop order factor $(1/p^{2})^{\frac{\epsilon}{2}}$ is implicitly included in the UV divergence result, and $L$-loop coupling factors introduce a $\mu^{\epsilon L}$ factor. Converting bare couplings $g_{0},h_{0}\to g,h$ introduces an overall $\left(\frac{\mu^{2}}{p^{2}}\right)^{\frac{\epsilon}{2}}$ factor, which contributes to the $\ln\frac{\mu^{2}}{p^{2}}$ expansion. Generally, $G_{\scriptsize\mbox{irre}}^{j{\tiny\mbox{-loop}}}$ carries an overall $\left(\frac{\mu^{2}}{p^{2}}\right)^{j\frac{\epsilon}{2}}$ factor, and we always treat the $(1/p^{2})^{\frac{\epsilon}{2}L}$ factor implicit in all UV divergence expressions.
The one-loop $\widetilde{Z}_{\varphi^{2}}$ is determined by enforcing UV finiteness of ${\widetilde{Z}_{\varphi^{2}}}G_{Q=2}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{Q})$, where we have used the fact that the 1PI contribution of $\widetilde{Z}_{\varphi^{1}}=1$. Rewriting bare couplings via eqn.(7) and expanding in the limit $g,h\to 0$ followed by $\epsilon\to 0$, the $1/\epsilon$ pole cancelation in the leading Log term gives,
$0=z_{\varphi^{2}}^{11}+g^{2}~~~\to~~~z_{\varphi^{2}}^{11}=-g^{2}~.$ (67)
No additional conditions arise from sub-Log terms. For the two-loop $\widetilde{Z}_{\varphi^{2}}$-factor, $G_{\scriptsize\mbox{irre}}^{{\tiny\mbox{2-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{2})$ is included. There are 5 independent irreducible topologies generating 7 cut graphs, with symmetry factors
@@
UV divergences are again computed by TwoPoint command in HyperlogProcedures, which are
@@
The two-loop irreducible UV contribution for $\varphi^{2}\to 2\varphi$ is
@@
Expanding ${\widetilde{Z}_{\varphi^{2}}}G_{Q=2}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{Q})$, the UV finiteness conditions, from canceling $1/\epsilon$ and $1/\epsilon^{2}$ poles in the leading Log term, are
@@
Using $z_{\varphi^{2}}^{11}$, solving these equations yields
@@
Notably, the sub-Log term’s $1/\epsilon$ pole cancelation provides an additional equation $0=g^{2}z_{\varphi^{2}}^{11}+g^{4}$, enabling full determination of $z_{\varphi^{2}}^{11},z_{\varphi^{2}}^{21},z_{\varphi^{2}}^{22}$ even without prior one-loop computations.
The $\widetilde{Z}_{\varphi^{3}}$-factor is determined via UV finiteness of $\frac{\widetilde{Z}_{\varphi^{3}}}{\widetilde{Z}_{\varphi^{2}}}G_{Q=3}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},\mbox{\textst{$x$}}_{2},x_{Q})$. The $\varphi^{3}$-$3\varphi$ correlation function topologies, as shown in Fig.(5),
@@
include irreducible topologies from $\varphi^{2}\to 2\varphi$, plus two new two-loop irreducible topologies $\Gamma^{\tiny\mbox{L2-XI}}_{\otimes\varphi{\varphi}\varphi}$ and $\Gamma^{\tiny\mbox{L2-XII}}_{\otimes\varphi{\varphi}\varphi}$. From eqn.(35), we get
@@
Since $G_{\scriptsize\mbox{irre}}^{{\tiny\mbox{1-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{2})$ and $G_{\scriptsize\mbox{irre}}^{{\tiny\mbox{2-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{2})$ have already been computed in $\varphi^{2}$ case as (66) and (77), only the new two-loop irreducible topologies require computation. These generate 4 cut graphs, with symmetry factors and coupling factors as,
@@
Then the required UV contribution of 2-loop irreducible topologies for $\varphi^{3}\to 3\varphi$ is
@@
Expanding $\frac{\widetilde{Z}_{\varphi^{3}}}{\widetilde{Z}_{\varphi^{2}}}G_{Q=3}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},\mbox{\textst{$x$}}_{2},x_{Q})$, the UV finiteness conditions, from canceling $1/\epsilon$ and $1/\epsilon^{2}$ poles in the leading Log term, and $1/\epsilon$ pole in the sub-Log term, give three equations,
@@
Solving these yields,
@@
With results (67), (80) and (90), the two-loop $\widetilde{Z}_{\varphi^{2}}$, $\widetilde{Z}_{\varphi^{3}}$-factors are finally determined as,
@@
The $\gamma_{Q}:=-\frac{\partial\ln\widetilde{Z}_{\varphi^{Q}}}{\partial\ln\mu}$ are computed using $\beta(g)$ and $\beta(h)$ from eqns.(61) and (62),
$\gamma_{2}=-\frac{\partial\ln\widetilde{Z}_{\varphi^{2}}}{\partial g}\beta(g)-\frac{\partial\ln\widetilde{Z}_{\varphi^{2}}}{\partial h}\beta(h)=-g^{2}+\left(\frac{7g^{4}N}{72}-\frac{49g^{4}}{36}-\frac{g^{3}h}{2}+\frac{7g^{2}h^{2}}{72}\right)~,$ (92)
$\gamma_{3}=-\frac{\partial\ln\widetilde{Z}_{\varphi^{3}}}{\partial g}\beta(g)-\frac{\partial\ln\widetilde{Z}_{\varphi^{3}}}{\partial h}\beta(h)=-3g^{2}+\left(\frac{7g^{4}N}{24}-\frac{67g^{4}}{12}-\frac{5g^{3}h}{2}+\frac{7g^{2}h^{2}}{24}\right)~.$ (93)
Assuming $\gamma_{Q}$ follows the ansatz $\gamma_{Q}=c_{1}Q(Q-1)(Q-2)+c_{2}Q(Q-1)$, we determine
@@
Then the scaling dimension of the $\varphi^{Q}$ operator is
$\Delta_{Q}=\left(\frac{d}{2}-1\right)Q+\gamma_{Q}+Q\gamma_{\varphi}~,$ (95)
where $\gamma_{\varphi}$ is the field anomalous dimension from eqn.(63).
@@ 4 Five-loop anomalous dimensions
The two-loop computation procedures outlined in the preceding section can be straightforwardly generalized to higher-loop calculations following an identical framework. The primary distinction lies in the need to account for a substantially larger number of Feynman diagrams, as topological complexity increases with loop order. The field renormalization factors $Z_{\varphi}$ and $Z_{\sigma}$ are determined from their respective two-point topologies, following the approach analogous to (41). Similarly, the $Z_{h}$ and $Z_{g}$-factors are derived from two-point integrals of cut graphs, generated by removing one external leg from the respective three-point topologies, following the approach outlined in (55). For the composite operator renormalization factors $\widetilde{Z}_{\varphi^{Q}}$, these are determined from two-point integrals of cut graphs produced by removing (Q1)(Q-1) external legs from irreducible 1PI $\varphi^{Q}\to Q\varphi$ topologies, in a manner parallel to that described in eqn.(35). The anomalous dimension $\gamma_{\varphi^{Q}}$ of the $\varphi^{Q}$ operator is then given by $\gamma_{Q}+Q\gamma_{\varphi}$, where $\gamma_{Q}$ is computed from $\widetilde{Z}_{\varphi^{Q}}$ and the beta functions. Determining the 5-loop $\gamma_{\varphi^{Q}}$ for general $Q$ requires results of $L$-loop $\gamma_{{Q_{L}}}$ with $Q_{L}=2,\ldots,L+1$ up to 5-loops. Notably, since UV divergences are evaluated from scalar Feynman integrals, the specific internal field configurations of the diagrams need not be considered, only their topological structures are relevant.
All topologies for two-point, three-point and $\varphi^{Q}$-$Q\varphi$ correlation functions, along with their corresponding cut graphs, are generated using custom Mathematica codes. The number of 1PI topologies for two-point, three-point functions and the irreducible correlation functions is listed in Table.(2).
@@
The number grows rapidly with increasing loop order. Cut graphs are generated via non-isomorphic enumeration of all possible external leg removals, and each cut graph must be individually evaluated to extract its UV divergence. As shown in the Table.(2), the number of cut graphs to be computed is approximately 3–4 times larger than the number of original topologies.
UV divergences are calculated using TwoPoint command in the HyperlogProcedures Maple package. The computational feasibility and time cost of these evaluations are highly sensitive to both the loop order and the truncation order of the dimensional regulator $\epsilon$. For an $L_{\tiny\mbox{max}}$-loop renormalization computation, each $L$-loop divergence in dimensional regularization must be computed up to the $\epsilon^{L_{\tiny\mbox{max}}-1-L}$ order. Increasing $L_{\tiny\mbox{max}}$ by one loop requires extending the $\epsilon$ expansion of all loop integrals by an additional order, which substantially increases computational time. For 5-loop computations, each $L$-loop divergence must be evaluated to the $\epsilon^{4-L}$ order, and the current version of HyperlogProcedures is capable of evaluating all required two-point integrals. Time consumption varies depending on the computer’s CPU performance and memory capacity. On a standard desktop workstation equipped with a 3.20GHz base frequency CPU and 64GB of RAM, the approximate time required to evaluate all necessary loop integrals is summarized in Table.(3).
@@
Time consumption also scales rapidly with increasing loop order. However, for 5-loop computations, it remains within a reasonably acceptable range. In contrast, the generation of symmetry factors and coupling factors for these topologies is nearly instantaneous, with negligible computational overhead compared to the time-intensive evaluation of UV divergences.
We have computed the field renormalization factors $Z_{\varphi}$, $Z_{\sigma}$, coupling renormalization factors $Z_{h}$, $Z_{g}$, and all composite operator renormalization factors $\widetilde{Z}_{\varphi^{Q_{L}}},Q_{L}=2,\ldots,L+1$ up to 5-loops. From these results, we have derived the beta functions $\beta(g),\beta(h)$, field anomalous dimensions $\gamma_{\varphi}$, $\gamma_{\sigma}$, and scaling dimensions of the $\varphi^{Q}$} operator, all up to 5-loops. These results are provided in ancillary files to the arXiv, and we have cross-checked them against existing literature with excellent agreements. For example, the up to 3-loop pieces of $\gamma_{\varphi},\gamma_{\sigma},\beta_{g},\beta_{h}$ are identical to those in Fei:2014xta , after unifying possible overall factors $(4\pi)$ from different conventions. The up to 4-loop pieces of results are identical to those in the $O(N)$ sector in Gracey:2015tta , after unifying possible overall factors $2$ from different conventions, and the fixed point solutions of beta functions up to $\epsilon^{4}$ order are identical to those in Kompaniets:2021hwg after compensating an overall $\sqrt{2}$ factor due to different convention.
The 5-loop correction to the anomalous dimension $\gamma_{Q}+Q\gamma_{\varphi}$ is presented in Appendix §A. The scaling dimension of the $\varphi^{Q}$ operator at the Wilson-Fisher fixed point is computed as
$\Delta_{Q\tiny\mbox{FP}}=\left(2-\frac{\epsilon}{2}\right)Q+\left(\gamma_{Q}+Q\gamma_{\varphi}\right)\Big|_{g=g^{\ast},h=h^{\ast}}~,$ (96)
where $g^{\ast},h^{\ast}$ are the fixed-point solutions of the beta functions. Our 5-loop fixed point solutions, provided in Appendix §B, are in perfect agreement with the literature Kompaniets:2021hwg . By substituting these fixed-point solutions into the anomalous dimensions yields $\Delta_{Q}$ in the large $N$ limit. The leading, sub-leading and sub-sub-leading order contributions of the large $N$ expansion are as follows,